OpenStudy (anonymous):

Find all rational zeros of the polynomial P(x)= 2 x^4 + 7 x^3 - 13 x^2 + 7 x - 15 x1 ≤ x2 ≤ x3 ≤ x4

Nnesha (nnesha):

\[\huge\rm \frac{ p }{ q } \]method to find rational zeros where p is factor of constant term and q is factor of leading coefficient

Nnesha (nnesha):

leading coefficient is 2 so i would replace q with factors of 2 \[\huge\rm \frac{ p }{ \pm1 , \pm 2 }\] write factors of constant term at the numerator :=)) leading coefficient is 2 and rational zero must be the divisor of constant term you should test these numbers to find zero replace x with each number into the equation if you get 0 then then u substituted would be the answer :=)

Nnesha (nnesha):

here is an example f(x)= 2x^3 +3x^2 +4 use the p/q expression \[\rm \frac{ factors ~of ~constant~term }{factors~of~leading~coefficeint }\] leading coefficient is 2 constant term 4 \[\frac{ \pm1 ,\pm 2 \pm 4 }{ \pm 1, \pm 2 }\] so possible rational zeros are  after finding possible rational zero i can test each number by plugging them into the original equation f(1)=2(1)^3+3(1)^2+4 f(2)=2(2)2+3(2)^2+4 same with other 3 options